A parking lot contains a total of 48 cars and
motorcycles. There are a total of 172 tires (not
counting spare tires) in the lot. Assuming each car has 4 tires and each motorcycle has 2 tires, determine
how many cars and how many motorcycles are in the parking lot.

Answer by ankor@dixie-net.com(22740)

A parking lot contains a total of 48 cars and motorcycles.
There are a total of 172 tires (not counting spare tires) in the lot
Assuming each car has 4 tires and each motorcycle has 2 tires, determine
how many cars and how many motorcycles are in the parking lot.
:
let c = no. of cars
let m = no. of motorcycles
:
"A parking lot contains a total of 48 cars and motorcycles."
c + m = 48
" There are a total of 172 tires"
4c + 2m = 172
multiply the 1st equation by 2, and subtract for the above equation
4c + 2m = 172
2c + 2m = 96
----------------subtraction eliminates m, find c
2c + 0 = 76
c = 76/2
c = 38 cars,
and
48-38 = 10 motorcycles
:
:
Check: 4(38) + 2(10) = 172